OpenStudy (anonymous):

i need help solving this. if f(x)=(6x^2)-x=3 and g(x)=3x=5 what is f(g(x)) and g(f(x))?

OpenStudy (anonymous):

\[f(g(x))=f(3x+5)=6(3x+5)^2-x-3\] i am assuming there is a typo in the question

OpenStudy (anonymous):

then algebra to get \[f(g(x))=6 x^2+35 x+51\]

OpenStudy (anonymous):

oh wait, maybe that is \[f(x)=6x^2-3+3\] and \[g(x)=3x+5\] not clear so no use continuing until it is clarified. answers will of course depend on what the equal sign should really be